How many milliliters of
11.5 M HCl(aq)
are needed to prepare
490.0 mL
of
1.00 M HCl(aq)
1
Expert's answer
2019-07-26T03:16:46-0400

"M1*V1=M2*V2"

"V1=M2*V2\/V1=490mL*1M\/11.5M=42.61mL"

Answer:42.61mL